# Learning Dense Correspondence for NeRF-Based Face Reenactment

### 16 Dec 2023 v1

@@ Abstract
Face reenactment is challenging due to the need to establish dense correspondence between various face representations for motion transfer. Recent studies have utilized Neural Radiance Field (NeRF) as fundamental representation, which further enhanced the performance of multi-view face reenactment in photo-realism and 3D consistency. However, establishing dense correspondence between different face NeRFs is non-trivial, because implicit representations lack ground-truth correspondence annotations like mesh-based 3D parametric models (e.g., 3DMM) with index-aligned vertexes. Although aligning 3DMM space with NeRF-based face representations can realize motion control, it is sub-optimal for their limited face-only modeling and low identity fidelity. Therefore, we are inspired to ask: Can we learn the dense correspondence between different NeRF-based face representations without a 3D parametric model prior? To address this challenge, we propose a novel framework, which adopts tri-planes as fundamental NeRF representation and decomposes face tri-planes into three components: canonical tri-planes, identity deformations, and motion. In terms of motion control, our key contribution is proposing a Plane Dictionary (PlaneDict) module, which efficiently maps the motion conditions to a linear weighted addition of learnable orthogonal plane bases. To the best of our knowledge, our framework is the first method that achieves one-shot multi-view face reenactment without a 3D parametric model prior. Extensive experiments demonstrate that we produce better results in fine-grained motion control and identity preservation than previous methods. Project page (video demo): https://songlin1998.github.io/planedict/.
Figure 1: Results of one-shot multi-view face reenactment. We present both cross-identity reenactment and multi-view synthesis at various yaw and pitch angles. Through comparisons with state-of-the-art HiDeNeRF, we illustrate that our PlaneDict module excels in fine-grained motion control, particularly for non-facial elements such as eyes, and offers better identity preservation (vertex distance from source identitynormal-$\downarrow$/ vertex distance from driving identitynormal-$\uparrow$) than utilizing 3DMM as correspondence.
@@ Introduction
$\mathcal{L}_{recons}=\|I_{d}-I\|_{1}+\lambda_{3}\|\mathcal{M}(I_{d})-\mathcal{M}(I)\|_{1},$ (7)
@@
The cross-identity reenactment is using the source and driving images of different identities, which is a more difficult challenge for source identity preservation and fine-grained motion transfer between different face tri-planes. We first qualitatively compare different state-of-the-art methods and show their synthesized results in Fig. 4. Although the fusion of identity and motion information is a hard problem, our framework with the PlaneDict module is able to generate cross-identity reenactment results with better image quality
@@
and identity fidelity without any artifacts. To further compare the identity preservation ability of our framework, as shown in Fig. 5, we evaluate and visualize the identity fidelity (i.e., the shape preservation of the source identity) using 3D face reconstruction models (Deng et al. 2019). The quantitative experiments are performed with VoxCeleb1 (Nagrani, Chung, and Zisserman 2017), VoxCeleb2 (Chung, Nagrani, and Zisserman 2018), and TalkingHead-1KH (Wang, Mallya, and Liu 2021), which are shown in Tab. 3 respectively.
@@
Ablation Study
As shown in Tab. 5, in the ablation study of whether to use the PlaneDict module, we adopt the same structure as identity deformations for obtaining motion and they are trained with the same time and dataset. Limited by the hardware, the max number of plane bases is 23. However, since 20, there has been almost no improvement. Therefore, we adopt 20 as
@@
the number of plane bases in our PlaneDict module to balance the quality and optimization difficulty.

### 19 Dec 2023 v2

@@ Abstract
Face reenactment is challenging due to the need to establish dense correspondence between various face representations for motion transfer. Recent studies have utilized Neural Radiance Field (NeRF) as fundamental representation, which further enhanced the performance of multi-view face reenactment in photo-realism and 3D consistency. However, establishing dense correspondence between different face NeRFs is non-trivial, because implicit representations lack ground-truth correspondence annotations like mesh-based 3D parametric models (e.g., 3DMM) with index-aligned vertexes. Although aligning 3DMM space with NeRF-based face representations can realize motion control, it is sub-optimal for their limited face-only modeling and low identity fidelity. Therefore, we are inspired to ask: Can we learn the dense correspondence between different NeRF-based face representations without a 3D parametric model prior? To address this challenge, we propose a novel framework, which adopts tri-planes as fundamental NeRF representation and decomposes face tri-planes into three components: canonical tri-planes, identity deformations, and motion. In terms of motion control, our key contribution is proposing a Plane Dictionary (PlaneDict) module, which efficiently maps the motion conditions to a linear weighted addition of learnable orthogonal plane bases. To the best of our knowledge, our framework is the first method that achieves one-shot multi-view face reenactment without a 3D parametric model prior. Extensive experiments demonstrate that we produce better results in fine-grained motion control and identity preservation than previous methods.
Figure 1: Results of one-shot multi-view face reenactment. We present both cross-identity reenactment and multi-view synthesis at various yaw and pitch angles. Through comparisons with state-of-the-art HiDeNeRF, we illustrate that our PlaneDict module excels in fine-grained motion control, particularly for non-facial elements such as eyes, and offers better identity preservation (vertex distance from source $\downarrow$/ vertex distance from driving $\uparrow$) than utilizing 3DMM as correspondence.
@@ Introduction
$\mathcal{L}_{recons}=\|\bm{I_{d}}-\bm{I}\|_{1}+\lambda_{3}\|\mathcal{M}(\bm{I_{d}})-\mathcal{M}(\bm{I})\|_{1},$ (7)
@@
The cross-identity reenactment is using the source and driving images of different identities, which is a more difficult challenge for source identity preservation and fine-grained motion transfer between different face tri-planes. We first qualitatively compare different state-of-the-art methods and
@@
show their synthesized results in Fig. 4. Although the fusion of identity and motion information is a hard problem, our framework with the PlaneDict module is able to generate cross-identity reenactment results with better image quality and identity fidelity without any artifacts. To further compare the identity preservation ability of our framework, as shown in Fig. 5, we evaluate and visualize the identity fidelity (i.e., the shape preservation of the source identity) using 3D face reconstruction models (Deng et al. 2019). The quantitative experiments are performed with VoxCeleb1 (Nagrani, Chung, and Zisserman 2017), VoxCeleb2 (Chung, Nagrani, and Zisserman 2018), and TalkingHead-1KH (Wang, Mallya, and Liu 2021), which are shown in Tab. 3 respectively.
@@
Ablation Study
As shown in Tab. 5, in the ablation study of whether to use the PlaneDict module, we adopt the same structure as identity deformations for obtaining motion and they are trained with the same time and dataset. Limited by the hardware, the max number of plane bases is 23. However, since 20, there has been almost no improvement. Therefore, we adopt 20 as the number of plane bases in our PlaneDict module to balance the quality and optimization difficulty.